OpenStudy (anonymous):

find the x intercept of the line 3x+8y=11

OpenStudy (lulubj):

\[\left( \frac{ 11 }{ 3 } ,0 \right)\]

jimthompson5910 (jim_thompson5910):

replace y with 0 3x+8y=11 3x+8*0=11 3x+0=11 3x = 11 now isolate x by dividing both sides by 3 3x=11 3x/3 = 11/3 x = 11/3 you can decide to leave it as a fraction or convert to decimal form

OpenStudy (deathlynine):

Yeah 11/3, usually on this question for 3x and 8y you would divide by 3.